# jina-embeddings-v4: Universal Embeddings for Multimodal Multilingual Retrieval

### v1

@@ Abstract
We introduce jina-embeddings-v4, a 3.8 billion parameter multimodal embedding model that unifies text and image representations through a novel architecture supporting both single-vector and multi-vector embeddings in the late interaction style. The model incorporates task-specific Low-Rank Adaptation (LoRA) adapters to optimize performance across diverse retrieval scenarios, including query-based information retrieval, cross-modal semantic similarity, and programming code search. Comprehensive evaluations demonstrate that jina-embeddings-v4 achieves state-of-the-art performance on both single-modal and cross-modal retrieval tasks, with particular strength in processing visually rich content such as tables, charts, diagrams, and mixed-media formats. To facilitate evaluation of this capability, we also introduce Jina-VDR, a novel benchmark specifically designed for visually rich image retrieval.
@@ 1 Introduction
This report presents jina-embeddings-v4, a multimodal embedding model capable of processing text and image data to produce semantic embedding vectors of varying lengths, optimized for a broad array of applications. It incorporates optimized LoRA adapters Hu et al. (2022) for information retrieval and semantic text similarity. An adapter is also provided for programming language embeddings, technical question-answering, and natural language code retrieval. It also brings new functionality to processing visually rich images (also called visual documents), i.e., materials mixing texts and images, containing tables, charts, diagrams, and other kinds of common mixed media Ding et al. (2024). We have also developed Jina-VDR, a new multilingual, multi-domain benchmark suite for a broad range of visual retrieval tasks, to evaluate the capabilities of jina-embeddings-v4.
This report discusses the challenges of developing a multimodal, multi-functional, state-of-the-art embedding model capable of handling texts in a variety of languages, including computer coding languages, images, and “visually rich” data. The resulting model, jina-embeddings-v4, projects inputs from all modalities into a unified semantic space, minimizing or eliminating the “modality gap” that has troubled similar projects Liang et al. (2022a). In addition, we introduce Jina-VDR, an advanced benchmark for images like screenshots and scans of visually complex documents.
@@
  • We have made particularly strong progress in handling visually rich images, especially for tasks outside of the existing ViDoRe benchmark Faysse et al. (2025), which is limited to OCR and question-answering. jina-embeddings-v4 outperforms other multimodal models by a significant margin on this type of material and supports a much more diverse set of use scenarios.
@@ 3 Related Work
Contrastive text-image training has led to ground-breaking results in zero-shot image classification and cross-modal retrieval in conjunction with dual encoder architectures like CLIP Radford et al. (2021). However, recent work shows better performance from vision-language models (VLMs) like Qwen2.5-VL Bai et al. (2025). Jiang et al. (2024) show that VLMs suffer less from a modality gap than dual encoder architectures.
@@
[…]
@@ 4 Model Architecture
Figure 1: Model Architecture
The architecture of jina-embeddings-v4, schematized in Figure 1, uses the Qwen2.5-VL model Bai et al. (2025) as a backbone. Like Qwen2.5-VL, it supports two input modalities, text and image, and processes its inputs in the same way. However, jina-embeddings-v4 has extended the architecture to support two output options: single-vector and multi-vector, and adds LoRA extensions to improve task-specific performance.
@@ 4.1 True Multimodal Processing
The Qwen2.5-VL paradigm differs from CLIP-style dual-encoder models in offering a single processing path that’s truly multimodal.
For text input, jina-embeddings-v4 and Qwen2.5-VL initially behave like other transformer-based embedding models: The text is tokenized, each token is replaced with a vector representation from a lookup table, and then these vectors are stacked and presented to a large language model (LLM).
@@
The Qwen2.5-VL paradigm used in jina-embeddings-v4 also includes a discrete image model but uses it in a different way. It produces a multi-vector result, comparable to late interaction models, and then passes this output to the LLM, as if it were a sequence of vectorized text tokens. The image embedding model acts as a preprocessor for the LLM, converting the image into what amounts to a sequence of vectorized “image tokens.”
@@ 4.2 Dual Mode Output
In contrast to Qwen2.5-VL and other embedding models in general, users can choose between two output options: Traditional single (dense) vector embeddings and ColBERT-style multi-vector embeddings for late interaction strategies.
@@ 5 Training Method
Before training, model weights are initialized to match Qwen/Qwen2.5-VL-3B-Instruct [^1] https://huggingface.co/Qwen/Qwen2.5-VL-3B-Instruct (Access: 06-04-2025). The multi-vector projection layer and LoRA adapters are randomly initialized. The weights of the backbone model are not modified during the training process. The LoRA adapters modify the effect of the backbone model layers and the projection layer. Only the adapters are trained.
@@ 5.1 Pair Training
We generate normalized single-vector and multi-vector embeddings for all texts and images in the selected pairs. We then construct a matrix of similarity values $\textbf{S}_{dense}(\mathcal{B})$ by calculating the cosine similarity of all combinations of single-vector embeddings $\bm{q}_{i}$ and $\bm{p}_{j}$ in $\mathcal{B}$. We construct an analogous matrix $\textbf{S}_{late}$ for each $\mathcal{B}$ for the multi-vector embeddings using a slightly modified version of Equation (1) to calculate their similarity. Our choice of loss function requires a normalized score, so we divide the late interaction score by the number of tokens in the query:
$s^{\prime}_{late}(q_{i},p_{j})=\frac{s_{late}(q_{i},p_{j})}{t}$ (2)
@@
[…]
@@
[…]
@@ 5.2.1 Asymmetric Retrieval Adapter
Our dataset of text hard negatives is similar to the data used to train jina-embeddings-v3 Sturua et al. (2024). We rely on existing datasets to create multimodal hard negatives for training, including Wiki-SS Ma et al. (2024) and VDR multilingual [^2] https://huggingface.co/datasets/llamaindex/vdr-multilingual-train (Access: 06-04-2025), but also mined hard negatives from curated multimodal datasets.
@@ 5.2.2 Text Matching Adapter
Symmetric semantic similarity tasks require different training from asymmetric retrieval. We find that training data with ground truth similarity values works best for this kind of task. As discussed in Sturua et al. (2024), we use the CoSENT [^3] https://github.com/bojone/CoSENT (Accessed: 08-28-2024) loss function $\mathcal{L}_{\mathrm{co}}$ from  Li and Li (2024):
@@ 5.2.3 Code Adapter
The backbone LLM Qwen2.5-VL was pre-trained on data including the StackExchangeQA [^4] https://github.com/laituan245/StackExchangeQA and the CodeSearchNet Husain et al. (2020) datasets, giving it some capacity to support code embeddings before further adaptation. Our LoRA training used the same triplet-based method described in Section 5.2.1. Training triplets are derived from a variety of sources, including CodeSearchNet, CodeFeedback Zheng et al. (2024), APPS Hendrycks et al. (2021), and the CornStack dataset Suresh et al. (2025).
@@ 7.1 Multilingual Text Retrieval
MTEB and MMTEB Enevoldsen et al. (2025) are the most widely used text retrieval benchmarks. For most tasks, we have used the asymmetric retrieval adapter, but for some symmetric retrieval tasks like ArguAna [^12] https://huggingface.co/datasets/mteb/arguana (Access: 06-21-2025). For evaluation, we prepend the query with the prefix “Given a claim, find documents that refute the claim” to reflect the task’s focus on retrieving passages that contradict (rather than support) the input claim, similar to  Wang et al. (2023)., we have used the text matching adapter instead. The results are tabulated in Appendix A.4.
For the MTEB benchmarks, which are all in English, see Table A11, for BEIR, see Table A12, and for the multilingual MMTEB, see Table A13. The performance of this new model is generally better than our previous model jina-embeddings-v3 and broadly comparable with the state-of-the-art.
@@ 7.2 Textual Semantic Similarity
We evaluated jina-embeddings-v4 with text-based semantic similarity (STS) benchmarks. The results for MTEB STS and MMTEB STS benchmarks are tabulated in Tables A15 and A16 of Appendix A.4 respectively. Our results are competitive with the state-of-the-art and are best-in-class for English similarity tasks.
@@ 7.3 Multimodal Retrieval
To evaluate the model’s performance on typical text-to-image search tasks, we used the common English and non-English tasks of the CLIP Benchmark [^13] https://github.com/LAION-AI/CLIP_benchmark (Access: 06-05-2025). The results are tabulated in Tables A8 to A10 of Appendix A.3. jina-embeddings-v4 has a higher average score than jina-clip-v2 and nllb-siglip-large, but the latter performs somewhat higher on the Crossmodal3600 benchmark Thapliyal et al. (2022) (see Table A10) because it includes content from low-resource languages not supported in jina-embeddings-v4’s Qwen2.5-VL backbone.
@@ A.4 MTEB and MMTEB
Table A12: Retrieval performance on MTEB BEIR (nDCG@10%).
Model Avg Arg CQA CF DB FEV FiQA HPQA MSM NFC NQ Quora SCD SCF TREC TOU
jina-embeddings-v3 53.91 54.33 42.36 42.36 41.00 89.06 47.35 64.67 40.82 36.62 64.23 89.09 19.81 72.31 77.72 26.30
jina-embeddings-v4 53.79 67.07 40.03 35.06 43.94 86.60 46.51 66.09 35.58 34.44 61.68 87.79 21.47 76.10 80.36 24.09
bge-m3 54.04 31.41 16.31 64.37 54.85
text-embedding-3-large 55.43 57.99 47.54 30.28 44.76 87.94 55.00 71.58 40.25 42.07 61.27 89.05 23.07 77.77 79.56 23.36
e5-mistral-7b-intr 57.07 61.65 46.00 38.35 48.90 87.84 56.81 75.72 43.06 38.58 63.54 89.61 16.32 76.42 87.03 26.27
Tasks: AvgMean nDCG@10% over all tasks, Arg: ArguAna, CQA: CQADupstackRetrieval, CF: ClimateFEVER, DB: DBPedia, FEV: FEVER, FiQA: FiQA2018, HPQA: HotpotQA, MSM: MSMARCO, NFC: NFCorpus, NQ: Natural Questions, Quora: QuoraRetrieval, SCD: SCIDOCS, SCF: SciFact, TREC: TRECCOVID, TOU: Touche2020

### v2

@@ Abstract
We introduce jina-embeddings-v4, a 3.8 billion parameter multimodal embedding model that unifies text and image representations through a novel architecture supporting both single-vector and multi-vector embeddings in the late interaction style. The model incorporates task-specific Low-Rank Adaptation (LoRA) adapters to optimize performance across diverse retrieval scenarios, including query-document retrieval, semantic text similarity, and code search. Comprehensive evaluations demonstrate that jina-embeddings-v4 achieves state-of-the-art performance on both single-modal and cross-modal retrieval tasks, with particular strength in processing visually rich content such as tables, charts, diagrams, and mixed-media formats. To facilitate evaluation of this capability, we also introduce Jina-VDR, a novel benchmark specifically designed for visually rich image retrieval.
**footnotetext: Equal contribution.
@@ 1 Introduction
We present jina-embeddings-v4, a multimodal embedding model capable of processing text and image data to produce semantic embedding vectors of varying lengths, optimized for a broad array of applications. It incorporates optimized LoRA adapters Hu et al. (2022) for information retrieval and semantic text similarity. An adapter is also provided for programming language embeddings, technical question-answering, and natural language code retrieval. It also brings new functionality to processing visually rich images (also called visual documents), i.e., materials mixing texts and images, containing tables, charts, diagrams, and other kinds of common mixed media Ding et al. (2024). We have also developed Jina-VDR, a new multilingual, multi-domain benchmark suite for a broad range of visual retrieval tasks, to evaluate the capabilities of jina-embeddings-v4.
We discuss the challenges of developing a multimodal, multi-functional, state-of-the-art embedding model capable of handling texts in a variety of languages, including computer coding languages, images, and “visually rich” data. The resulting model, jina-embeddings-v4, projects inputs from all modalities into a unified semantic space, minimizing or eliminating the “modality gap” that has troubled similar projects Liang et al. (2022a). In addition, we introduce Jina-VDR, an advanced benchmark for images like screenshots and scans of visually complex documents.
@@
  • We have made particularly strong progress in handling visually rich images, especially for tasks outside of the existing ViDoRe benchmark Faysse et al. (2025), which is limited to question-answering. jina-embeddings-v4 outperforms other multimodal models by a significant margin on this type of material and supports a much more diverse set of use scenarios.
@@ 3 Related Work
Contrastive text-image training has led to ground-breaking results in zero-shot image classification and cross-modal retrieval in conjunction with dual encoder architectures like CLIP Radford et al. (2021). However, recent work shows better performance from vision-language models (VLMs) like Qwen2.5-VL-3B-Instruct Bai et al. (2025). Jiang et al. (2024) show that VLMs suffer less from a modality gap than dual encoder architectures.
@@
Similarity scoring in late interaction models does not use simple cosine similarity. Khattab and Zaharia (2020) Instead, similarity is calculated asymmetrically over two sequences of token embeddings — a query and a document — by summing up the maximum cosine similarity values of each query token embedding to any of the token embeddings from the document. Thus, for query embedding $q$ and document embedding $p$, their late interaction similarity score $s_{\mathrm{late}}(q,p)$ is determined by:
$s_{\mathrm{late}}(q,p)=\sum_{i=1}^{n}\max_{j\in\{1,\dots,m\}}\bm{q}_{i}\cdot\bm{p_{j}}^{T}$ (1)
@@ 4 Model Architecture
Figure 1: Architecture of jina-embeddings-v4. The model employs a unified LM built on the Qwen2.5-VL-3B-Instruct backbone (3.8B parameters). Text and image inputs are processed through a shared pathway: images are first converted to token sequences via a vision encoder, then both modalities are jointly processed by the language model decoder with contextual attention layers. Three task-specific LoRA adapters (60M parameters each) provide specialized optimization for retrieval, text-matching, and code search tasks without modifying the frozen backbone weights. The architecture supports dual output modes: (1) single-vector embeddings (2048 dimensions, truncatable to 128) generated via mean pooling for efficient similarity search, and (2) multi-vector embeddings (128 dimensions per token) via projection layers for the late interaction style retrieval.
The architecture of jina-embeddings-v4, schematized in Figure 1, employs a unified multimodal language model built on the Qwen2.5-VL-3B-Instruct [^1] https://huggingface.co/Qwen/Qwen2.5-VL-3B-Instruct backbone Bai et al. (2025). Text and image inputs are processed through a shared pathway: Images are first converted to token sequences via a vision encoder, then both modalities are jointly processed by the language model decoder with contextual attention layers. This unified design eliminates the modality gap present in dual-encoder architectures while maintaining competitive performance across text, image, and cross-modal tasks.
As shown in Figure 1, this architecture supports dual output modes, as outlined in Section 4.2. Furthermore, three task-specific LoRA adapters, each with 60M parameters, provide specialized task optimization without modifying the frozen backbone weights. These are described in Section 4.3.
@@ 4.1 True Multimodal Processing
The Qwen2.5-VL-3B-Instruct paradigm differs from CLIP-style dual-encoder models in offering a single processing path that’s truly multimodal.
For text input, jina-embeddings-v4 and Qwen2.5-VL-3B-Instruct initially behave like other transformer-based embedding models: The text is tokenized, each token is replaced with a vector representation from a lookup table, and then these vectors are stacked and presented to a large language model (LLM).
@@
The Qwen2.5-VL-3B-Instruct paradigm used in jina-embeddings-v4 also includes a discrete image model but uses it in a different way. It produces a multi-vector result, comparable to late interaction models, and then passes this output to the LLM, as if it were a sequence of vectorized text tokens. The image embedding model acts as a preprocessor for the LLM, converting the image into what amounts to a sequence of vectorized “image tokens.”
@@ 4.2 Dual Mode Output
In contrast to Qwen2.5-VL-3B-Instruct and other embedding models in general, users can choose between two output options: Traditional single (dense) vector embeddings and ColBERT-style multi-vector embeddings for late interaction strategies.
@@ 5 Training Method
Before training, model weights are initialized to match Qwen/Qwen2.5-VL-3B-Instruct. The multi-vector projection layer and LoRA adapters are randomly initialized. The weights of the backbone model are not modified during the training process. The LoRA adapters modify the effect of the backbone model layers and the projection layer. Only the adapters are trained.
@@ 5.1 Pair Training
We generate normalized single-vector and multi-vector embeddings for all texts and images in the selected pairs. We then construct a matrix of similarity values $\textbf{S}_{\mathrm{dense}}(\mathcal{B})$ by calculating the cosine similarity of all combinations of single-vector embeddings $\bm{q}_{i}$ and $\bm{p}_{j}$ in $\mathcal{B}$. We construct an analogous matrix $\textbf{S}_{\mathrm{late}}$ for each $\mathcal{B}$ for the multi-vector embeddings using a slightly modified version of Equation (1) to calculate their similarity. Our choice of loss function requires a normalized score, so we divide the late interaction score by the number of tokens in the query:
$s^{\prime}_{\mathrm{late}}(q_{i},p_{j})=\frac{s_{\mathrm{late}}(q_{i},p_{j})}{t}$ (2)
@@
$\mathcal{L}_{D}(B,\tau)$ $:=D_{\mathrm{KL}}(\textbf{S}^{\prime}_{\mathrm{dense}}(\mathcal{B})\,\|\,\textbf{S}^{\prime}_{\mathrm{late}}(\mathcal{B}))$
@@
$w_{1}\mathcal{L}_{\mathrm{NCE}}(\textbf{S}_{\mathrm{dense}}(\mathcal{B}_{\mathrm{txt}}),\tau)$
$+$ $w_{2}\mathcal{L}_{\mathrm{NCE}}(\textbf{S}_{\mathrm{late}}(\mathcal{B}_{\mathrm{txt}}),\tau)+w_{3}\mathcal{L}_{D}(\mathcal{B}_{\mathrm{txt}})$
$+$ $w_{4}\mathcal{L}_{\mathrm{NCE}}(\textbf{S}_{\mathrm{dense}}(\mathcal{B}_{\mathrm{multi}}),\tau)$
$+$ $w_{5}\mathcal{L}_{\mathrm{NCE}}(\textbf{S}_{\mathrm{late}}(\mathcal{B}_{\mathrm{multi}}),\tau)+w_{6}\mathcal{L}_{D}(\mathcal{B}_{\mathrm{multi}})$ (6)
@@ 5.2.1 Asymmetric Retrieval Adapter
Our dataset of text hard negatives is similar to the data used to train jina-embeddings-v3 Sturua et al. (2024). We rely on existing datasets to create multimodal hard negatives for training, including Wiki-SS Ma et al. (2024) and VDR multilingual [^2] https://huggingface.co/datasets/llamaindex/vdr-multilingual-train, but also mined hard negatives from curated multimodal datasets.
@@ 5.2.2 Text Matching Adapter
Symmetric semantic similarity tasks require different training from asymmetric retrieval. We find that training data with ground truth similarity values works best for this kind of task. As discussed in Sturua et al. (2024), we use the CoSENT [^3] https://github.com/bojone/CoSENT loss function $\mathcal{L}_{\mathrm{co}}$ from  Li and Li (2024):
@@ 5.2.3 Code Adapter
The backbone LLM Qwen2.5-VL-3B-Instruct was pre-trained on data including the StackExchangeQA [^4] https://github.com/laituan245/StackExchangeQA and the CodeSearchNet Husain et al. (2020) datasets, giving it some capacity to support code embeddings before further adaptation. Our LoRA training used the same triplet-based method described in Section 5.2.1. Training triplets are derived from a variety of sources, including CodeSearchNet, CodeFeedback Zheng et al. (2024), APPS Hendrycks et al. (2021), and the CornStack dataset Suresh et al. (2025).
@@ 7.1 Multilingual Text Retrieval
MTEB and MMTEB Enevoldsen et al. (2025) are the most widely used text retrieval benchmarks. For most tasks, we have used the asymmetric retrieval adapter, but for some symmetric retrieval tasks like ArguAna [^12] https://huggingface.co/datasets/mteb/arguana, we have used the text matching adapter instead. For evaluation, we prepend the query with the prefix “Given a claim, find documents that refute the claim” to reflect the task’s focus on retrieving passages that contradict, rather than support, the input claim, similar to  Wang et al. (2023). The results are tabulated in Appendix A.4.
For the MTEB benchmarks, which are all in English, see Table A11, and for the multilingual MMTEB, see Table A12. The performance of this new model is generally better than our previous model jina-embeddings-v3 and broadly comparable with the state-of-the-art.
@@ 7.2 Textual Semantic Similarity
We evaluated jina-embeddings-v4 with text-based semantic similarity (STS) benchmarks. The results for MTEB STS and MMTEB STS benchmarks are tabulated in Tables A14 and A15 of Appendix A.4 respectively. Our results are competitive with the state-of-the-art and are best-in-class for English similarity tasks.
@@ 7.3 Multimodal Retrieval
To evaluate the model’s performance on typical text-to-image search tasks, we used the common English and non-English tasks of the CLIP Benchmark [^13] https://github.com/LAION-AI/CLIP_benchmark. The results are tabulated in Tables A8 to A10 of Appendix A.3. jina-embeddings-v4 has a higher average score than jina-clip-v2 and nllb-siglip-large, but the latter performs somewhat higher on the Crossmodal3600 benchmark Thapliyal et al. (2022) (see Table A10) because it includes content from low-resource languages not supported in jina-embeddings-v4’s Qwen2.5-VL-3B-Instruct backbone.